Circular No. 6519

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Postal Address: Central Bureau for Astronomical Telegrams Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A. IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions)

BMARSDEN@CFA.HARVARD.EDU or DGREEN@CFA.HARVARD.EDU (science) Phone 617-495-7244/7440/7444 (for emergency use only)

COMET C/1995 O1 (HALE-BOPP) C. M. Wright, Max-Planck-Institut für Extraterrestrische Physik; and

C. H. Smith and T. Fujiyoshi, University College, Australian Defence Force Academy, report observations on Aug. 2.5 and 3.4 UT with the Anglo-Australian Telescope and the mid-infrared camera NIMPOL: “Although conditions were not photometric, estimated fluxes in a 5′′-diameter aperture on Aug. 2.5 at 8.5, 11.5 and 12.5 μm (1-μm-wide filters) were (in units of 10^{−18} W cm^{−2} μm^{−1}), 4.4 ± 0.2, 11.7 ± 0.1 and 11.7 ± 0.2, respectively. Quoted uncertainties are purely statistical, whilst absolute calibration is ∼ 30 percent. The 8.5–12.5-μm color temperature is 187±10 K, well above the 147 K expected from a blackbody in radiative equilibrium at 3.57 AU from the sun. The silicate feature, as judged from the 11.5-μm data, is a factor of 1.1 ± 0.2 above the 187-K continuum. On Aug. 3.4 the 11.5-μm flux in a 5′′-diameter aperture was (in units as before) 6.6 ± 0.1, suggesting that the comet faded by at least 15 percent between the observations.”

Naked-eye magnitude estimates: Dec. 2.06 UT, 4.1 (A. Hale, Cloudcroft, NM); 3.69, 3.8 (K. Hornoch, Lelekovice, Czech Republic); 4.06, 4.0 (Hale); 9.74, 4.2 (S. Garro, Orcières-Merlette, France).

XTE J1856+053 AND GRO J1849–03 D. Barret, J. E. Grindlay and P. F. Bloser, Harvard-Smithsonian Cen-

ter for Astrophysics (CfA); and B. A. Harmon, S. N. Zhang, C. A. Wilson, C. R. Robinson and W. S. Paciesas, NASA Marshall Space Flight Center, report for the BATSE Team: “As part of the galactic-plane survey for faint x-ray transients (Grindlay et al. 1996, A.Ap. Suppl., in press), the CfA BATSE Image Search (CBIS) system for automated analysis of earth-occultation images (Barret et al. 1996, in preparation) was used to scan sky images of the Aql-Sct-Ser region. In a 20–50-keV image integrated over Sept. 3–17 we found evidence of emission from the recently-discovered x-ray transient XTE J1856+053 (IAUC 6504). Also, in images made for the Sept. 3–17 and 16–30 time periods, CBIS reported weak emission from the recurrent transient GRO J1849–03, consistent with the 241-day period reported by Zhang et al. (IAUC 6150). Analysis of the lightcurves for XTE J1856+053 and GRO J1849-03 shows peak fluxes in the 20–100-keV band between 30 and 60 mCrab, with XTE J1856+053 peaking in BATSE data on Sept. 7–9, preceding the ASM peak flux by ∼ 8 days. This behavior is characteristic of low-mass-binary transients and could indicate the presence of a black hole. GRO J1849–03 has a very broad peak around Sept. 19.”

1996 December 9  \begin{matrix}c \\ ◯ \end{matrix} Copyright 1996 CBAT  Brian G. Marsden